# From Prompts to Printable Models: Support-Effective 3D Generation via Offset Direct Preference Optimization

### v1

@@ IV-D Discussions
Our approach is a preliminary effort to fine-tune a diffusion-based 3D generative model for specific applications, though some limitations remain. The reliance on the manifold assumption for simulating support structures may introduce biases, especially when dealing with double-sided meshes or floating facets. These issues are not inherent to our method and could be addressed with future geometry-assured 3D generation models. While SEG primarily targets FDM printers with a standard self-support angle, it is flexible and can be extended to other 3D printing technologies like SLA and SLS. Future work will explore optimizing upright orientation and self-support angles to enhance support efficiency across different printing methods.

### v2

@@ IV-D Discussions
Our approach is a preliminary effort to fine-tune a diffusion-based 3D generative model for specific applications, though some limitations remain. The reliance on the manifold assumption for simulating support structures may introduce biases, especially when dealing with double-sided meshes or floating facets. These issues are not inherent to our method and could be addressed with future geometry-assured 3D generation models. The problem of modifying existing designs to be largely orthogonal to our generative workflow, where both can serve overlapping user groups. Our method could be extended to other generative settings in the future. The training cost of our method follows the “train once, run everywhere” nature of foundation models. Each inference takes only  50 seconds on our single A100 GPU. Since our SEG is fine-tuned from a pretrained 3D generative model, some failure cases arise from limitations inherited from the underlying model. In such cases, it is difficult to clearly attribute the failures specifically to our attempt to minimize support structures. While SEG primarily targets FDM printers with a standard self-support angle, it is flexible and can be extended to other 3D printing technologies like SLA and SLS. Future work will explore optimizing upright orientation and self-support angles to enhance support efficiency across different printing methods.
@@ V Conclusion
Acknowledgment
The authors would thank the valuable help offered by the following students working in the Computational Robotics and Manufacturing Lab at the Chinese University of Hong Kong lead by Prof. Guoxin Fang, including Zhuo Huang and Richard Yixuan Zhou.